If 85.0 L of natural gas, which is essentially methane (CH4), undergoes complete combustion at STP, how many Liters of CO2 is produced

                 CH4 +  O2     →  CO2+  H2O  



1
Expert's answer
2021-03-24T01:11:07-0400

From the equation 1 mole of CH4 produces 1 mole of CO2

Molar Mass of CH4 = 16.04 g/mol

= 85 L = 85000 g / 16.04

= 5299.252 moles

Molar Mass of CO2 = 44.01g/mol

= 44.01 × 5299.252 = 233220.1

= 233220.1/1000

= 233.2201 L of CO2